Companies U and L are identical in every respect except that U is unlevered while L has $7 million of 6% bonds outstanding. Assume that (1) there are no corporate or personal taxes, (2) all of the other MM assumptions are met, (3) EBIT is $1 million, and (4) the cost of equity to Company U is 10%. What value would MM estimate for each firm? Enter your answers in millions. For example, an answer of $1.2 million should be entered as 1.2, not 1,200,000. Round your answers to two decimal places. Company U $ million Company L $ million What is rs for Firm U? Round your answer to two decimal places. % What is rs for Firm L? Do not round intermediate calculations. Round your answer to two decimal places. % Find SL. Round your answer to two decimal places. Enter your answer in millions. For example, an answer of $1.2 million should be entered as 1.2, not 1,200,000. $ million What is the WACC for Firm U? Round your answer to one decimal place. % What the WACC for Firm L? Round your answer to one decimal place. % Suppose VU = $20 million and VL = $22 million. According to MM, are these values consistent with equilibrium?
Answer #1

Company U

Company L

EBIT= $1 million

EBIT= $1 million

Cost of equity = 10%

-

6% Debt = $7 mil

No taxes

No taxes

  1. Value of unlevered firm with no tax = EBIT/ Cost of equity

Value of firm for company U= 1000000/10%

= $10,000,000 or $10 million

Value of firm of levered firm = value of unlevered firm + tax rate* value of debt

But for no tax firm , value of levered firm = value of unlevered firm

So value of L= $10 millions

    In case of unlevered firm R0 ( cost of equity)= Rs

    So Rs for company U= 10%

    For levered firm:

    ‘rs (return on equity)= (cost of Equity of unlevered firm*Equity/ Value of firm) + [(Cost of equity of unlevered firm-Cost of debt)*debt/ value of firm]

    Rs for company L= 0.1*3/10 + (0.1-0.06)*7/10

    = 0.03 + 0.028

    = 0.058

    = 5.8%

    So Rs of U=0.1

    Rs of L= 0.058

    1. S(L) = equity portion of the firm

    Market value of Equity = total value of firm- debt

    Here for company L

    S(L)= 10-7

    = $3 million

    1. WACC= {kd * debt/ debt+ equity}+ {ke*equity/debt+ equity}

    For U WACC= cost of equity = 10%

    For L WACC= 6% * 7/10) + 10%* 3/10

    = 0.042 +0.03

    = 0.072

    = 7.2%

    1. VU = $20 million

    VL= $22 million

    This shows that Vu is not equal to VL . and when there are no taxes, VU=VL

    So they are not at equilibrium.